OpenStudy (anonymous):

Which is the 10th term in the sequence:62, 47, 32, 17, 2, …?

OpenStudy (anonymous):

Well first how do you get from 62 to 47?

OpenStudy (anonymous):

62 - 47 = 15. So subtract 15.

OpenStudy (anonymous):

32 - 15 = 17 17 - 15 = 2 2 - 15 = ?

OpenStudy (anonymous):

no when you subtract you get a smaller number than the original

OpenStudy (anonymous):

Excuse me, i meant 73

OpenStudy (anonymous):

2 - 15 = -13 which is the 6th term -13 - 15 = -28 which is the 7th term -28 - 15 = -43 which is the 8th term -43 - 15 = -58 which is the 9th term -58 - 15 = -73 which is the 10th term

OpenStudy (anonymous):

-73 NOT 73; there is a huge difference between the two. The pattern started out from positive great numbers and then ends out with negative lower numbers because the pattern is decreasing by 15. When something decreases, it becomes lower eventually